Home‎ > ‎Reports‎ > ‎Crew098‎ > ‎

day04


Crew 98 Log Book for January 18, 2011

Photos of the day





Commander's Check-In Report

Virgiliu Pop Reporting

Crew Physical Status: Excellent

Time Departed/Returned from EVA: 3:50 PM-6:00 PM

Brief Narrative of Field Mission Results: Science team took a soil sample and a rock sample.

EVA Data/Interpretations: Samples to be analyzed.

Engineering/Hab Maintenance: As per the forthcoming engineering report.

Report Transmission Schedule: Intended by 8 PM.

Plans for Tomorrow: A new EVA to be performed.

Inventory: Apparently we are missing lots of pre-cooked (to-be-hydrated) meals, and some soldering materials.

Support Requested: There is still not enough grey water to be pumped in the toilet.

Miscellaneous: First sim day.


Commander's Report

Virgiliu Pop Reporting

First simulation day!

It started with a nice oatmeal breakfast, then we performed a thorough cleaning of the whole hab ... we decided that the lower deck is "slippers only', the upper deck is 'socks only', and the EVA dress room is 'boots only'. It is not true that men are from Mars and women are from Venus - they both belong to Mars, and a crew like our own, where women are the majority, is likely to keep Mars very clean.

We then had lunch - a whole eight portions of soup for a very tired crew of 5, while the mission commander got a nickname - Captain Cook! It is true, as a personal note, that I love to cook - and hate to clean up! - and I think the best way of managing a mission is to have crew members do what they like best (but also what they have to do ...).

After lunch, each member tended to their duties: Hari went to the Greenhab and later to the Musk observatory, Dragos drove the pressurized rover (SUV) with Betty and Iulia to the EVA site, while the commander had to set-up the rover. What a joy to do some electronics after more than 20 years - and what a chore to wash the dishes!

After the away team returned from the EVA, the crew wrote their reports, and we all look forward to a nice dinner and to our first astronomy lesson at the Musk observatory - courtesy of Haritina.

Engineering Report

Haritina Mogosanu Reporting

Generator/Electricity:


Kitty (Diesel Generator) (Used? Yes/No): used

Honey (Gas Generator) (Used? Yes/No): not used

Battery Charge Level:?
29.6 V everstart
34.5 V armor plate
40.6 V green

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):
1 generator run all day ok


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.5

Propane Tank (Full/0.75/0.5/0.25/Empty): 0.397

Gasoline Tank (Full/0.75/0.5/0.25/Empty):1 containers full x 5 gallons 2 half

Notes/Comments:


Oil: not checked DG said to leave it to him

Kitty Oil Quantity (# of Quarts in storage):2.5 gal

Honey Oil Quantity (# of Quarts in storage):1 gal

ATV Oil Quantity (# of Quarts in storage):on the engineering work bench
2 x 946 ml

Car Oil Quantity (# of Quarts in storage):where is it?


Water Status:

Outside Potable Water Tank Level (inches from bottom): 9

Trailer Potable Water Tank Level (inches from bottom): very low - little water left

Hab Potable Water Tank Level (inches from bottom): 14

Potable Water Meter Reading: 410363

Notes/Comments:


GreenHab:

Condition of plants in Tank 1: plants dead please advise

Condition of plants in Tank 2: very poor only 4 plants left

Condition of plants in Tank 3: improving

Greenhab Notes/Comments: plants are very few and have some leaves dead. I started the recovery procedure, removed dead leaves and separated some stolons. Plants look good now.
Tank 3 has water hyacinth, duckweed, some fresh water algae and one more plant that looks like moss and has a tiny root.
There is a Sagittaria sp in the second tank that started growth and has few tiny leaves. I removed all dead parts.


Observatory:

Telescope Used? (Yes/no) yes

Observatory Notes: telescope is severely out if alignment. I started aligning, first the finderscope. The Moon looks beautiful through it tonight.


Transportation:

Hab Car (NOT RENTAL)
Used (yes/no):Yes
Oil Checked (yes/no):No
Tire Status:not checked
Odometer at end of day: 123186
Notes/Comments on Hab Car:

ATV 1 (Opportunity):
Used: y/n n
Oil Checked: y/n n
Fuel Consumed: filled
Tire Status:not checked
Comments: started

ATV 2 (Spirit):n
Used: y/n y
Oil Checked: y/n n
Fuel Consumed:replaced
Tire Status:not checked
Comments: started

ATV 3 (Viking I):
Used: y/n n
Oil Checked: y/n n
Fuel Consumed: replaced
Tire Status: not checked
Comments: started

ATV 4 (Viking II):
Used: y/n y
Oil Checked: y/n n
Fuel Consumed: replaced
Tire Status:not checked
Comments:


Heating and Ventilation:
Thermostat setting upstairs (Farenheit): 58
Actual temperature upstairs (Farenheit): 79

Thermostat setting downstairs (Farenheit): 55
Actual temperature downstairs (Farenheit): 68


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: all ok


General Engineering Notes/Comments:

Science Report (Biology)

Beatrice Gilea Reporting

Time:16:00 - 18:00

Crew Members: Iulia Jivanescu, Dragos Bratasanu, Beatrice Gilea

Site Location:
1. N 38deg 24min 0,2sec
W 110deg 47min 27sec
Elevation 1359m
2. N 38deg 24min 0,4sec
W 110deg 47min 30sec
Elevation 1363m

Transit Mode: Pressurized rover (SUV)

Objectives:
Take rock and soil samples and bring them to the Hab for microbiological analysis

Results:
Samples not yet analyzed

Lessons Learned:
It is very difficult to take samples (to keep the instruments sterile) for microbiological analysis while wearing the EVA suit.

Science Report (Geology)

Iulia Jivanescu Reporting

Time: 16.00-18.00

Crew Members: Beatrice Gilea, Dragos Bratasanu, Iulia Jivanescu

Site Location: N 38deg 24 min 02 sec
W 110 deg 47 min 27 sec
N 38deg 24min 04sec
W 110deg 47min 30sec

Transit Mode: Pressurized rover (SUV)

Objectives: take rock and soil samples and analyze them for pH and radioactivity; sample them for the Biologist. collected rock samples in order to assess them on the Geological scale

Results: samples were not yet analysed

Lessons Learned: The use of the hand drill was much more difficult while wearing the EVA suit.

EVA 1 Report

Beatrice Gilea Reporting

Time: 16:00 - 18:00

Crew Members: Iulia Jivanescu, Dragos Bratasanu, Beatrice Gilea

Site Location:
1. N 38deg 24min 0,2sec
W 110deg 47min 27sec
Elevation 1359m
2. N 38deg 24min 0,4sec
W 110deg 47min 30sec
Elevation 1363m

Transit Mode: Pressurized rover (SUV)

Objectives: Rock and soil sampling

Results: Samples not yet analyzed

Lessons Learned: A lot more difficult to work in a EVA suit - dexterity and mobility are impaired. We will have to get used to using the radio for communication at all times.


Cook's Report

Virgiliu Pop Reporting

Spreadsheet of Cook's Report.